Reduce the quadratic form 8x²— 4xy + 5y² to
its normal canonical form.
1
Expert's answer
2020-03-25T13:06:54-0400

Reduce "8x^2-4xy+5y^2" to its normal canonical form.

It is possible to solve this problem using orthogonal transformation.

Let it be a matrix "A" such that

"A = \begin{bmatrix} \mathrm{coef} x_1^2 & \frac{1}{2} \mathrm{coef} x_1x_2 \\
\frac{1}{2}\mathrm{coef} x_2x_1 & \mathrm{coef} x_2^2
\end{bmatrix}"

Remembering that "x_1x_2=x_2x_1" it is possible to calculate "A":

"A = \begin{bmatrix} 8 & -2 \\
-2 & 5
\end{bmatrix}"

To express the given quadratic form in canonical form it is necessary to diagonalize matrix "A", that can be done by the following formula:

"M^{-1}AM =D",

where "M" is normilized eigenvectors matrix and "D" is diagonalized matrix "A".

To find "M" it is necessary to find eigenvalues and eigenvectors of matrix "A":

"\det(\lambda I-A)=0 \Leftrightarrow \begin{vmatrix}
\lambda-8 & 2 \\
2 & \lambda-5
\end{vmatrix} = (\lambda-8)(\lambda-5)-4=\lambda^2-13+36=0"

This equation can be solved with usage of Vieta's formulas:

"\begin{cases}
\lambda_1+\lambda_2=13 \\
\lambda_1 \cdot \lambda_2 = 36
\end{cases} \Bigg| \Rightarrow
\begin{cases}
\lambda_1=4 \\
\lambda_2=9
\end{cases}"

Eigenvector, corresponding to eigenvalue "\lambda_i" can be found with solving the following equation:

"Av_i=\lambda_iv_i",

where "v_i" is eigenvector.

1. Eigenvector, corresponding to eigenvalue "\lambda_1=4":

"Av_1=\lambda_iv_1 \Leftrightarrow \begin{cases}
8v_{11}-2v_{12}=4v_{11}\\
-2v_{11}+5v_{12}=4v_{12}
\end{cases} \Bigg| \Leftrightarrow \begin{cases}
v_{12}=2v_{11}\\
-2v_{11}+10v_{11}=8v_{11}
\end{cases} \Bigg|
\Leftrightarrow \begin{cases}
v_{12}=2v_{11}\\
8v_{11}=8v_{11}
\end{cases} \Bigg|
\Leftrightarrow \begin{cases}
v_{12}=2v_{11}\\
v_{11} \mathrm{-any\ number}
\end{cases}"

Let "v_{11}=1" , so "v_{12} = 2 v_{11} = 2", so normilized vector "v_1" can be found as:

"\overline{v}_1 = \begin{bmatrix}
\frac{v_{11}}{\sqrt{v_{11}^2+v_{12}^2}} \\
\frac{v_{12}}{\sqrt{v_{11}^2+v_{12}^2}}
\end{bmatrix} = \begin{bmatrix}
\frac{1}{\sqrt{1^2+2^2}} \\
\frac{2}{\sqrt{1^2+2^2}}
\end{bmatrix} =\begin{bmatrix}
\frac{1}{\sqrt{5}} \\
\frac{2}{\sqrt{5}}
\end{bmatrix}"

2. Eigenvector, corresponding to eigenvalue "\lambda_2=9":

"Av_2=\lambda_iv_2 \Leftrightarrow \begin{cases}
8v_{21}-2v_{22}=9v_{21}\\
-2v_{21}+5v_{22}=9v_{12}
\end{cases} \Bigg| \Leftrightarrow \begin{cases}
v_{22}=-\frac{1}{2}v_{21}\\
-2v_{21}-\frac{5}{2}v_{21}=-\frac{9}{2}v_{21}
\end{cases} \Bigg|
\Leftrightarrow \begin{cases}
v_{22}=-\frac{1}{2}v_{21}\\
-\frac{9}{2}v_{21}=-\frac{9}{2}v_{21}
\end{cases} \Bigg|
\Leftrightarrow \begin{cases}
v_{22}=-\frac{1}{2}v_{21}\\
v_{21} \mathrm{-any\ number}
\end{cases}"

Let "v_{21}=1" , so "v_{22} = -\frac{1}{2}v_{21} = -\frac{1}{2}", so normilized vector "v_2" can be found as:

"\overline{v}_2 = \begin{bmatrix}
\frac{v_{21}}{\sqrt{v_{21}^2+v_{22}^2}} \\
\frac{v_{22}}{\sqrt{v_{21}^2+v_{22}^2}}
\end{bmatrix} = \begin{bmatrix}
\frac{1}{\sqrt{1^2+(-\frac{1}{2})^2}} \\
\frac{-\frac{1}{2}}{\sqrt{1^2+(-\frac{1}{2})^2}}
\end{bmatrix} =\begin{bmatrix}
\frac{2}{\sqrt{5}} \\
-\frac{1}{\sqrt{5}}
\end{bmatrix}"

Combining "\overline{v}_1" and "\overline{v}_2" it is possible to get "M":

"M = \begin{bmatrix}
\overline{v}_1 & \overline{v}_2
\end{bmatrix} = \begin{bmatrix}
\frac{1}{\sqrt{5}} & \frac{2}{\sqrt{5}} \\
\frac{2}{\sqrt{5}} & -\frac{1}{\sqrt{5}}
\end{bmatrix} = \frac{1}{\sqrt{5}} \cdot \begin{bmatrix}
1 & 2 \\
2 & -1
\end{bmatrix}"

Inverse matrix "M^{-1}" can be found as:

"M^{-1} = \cfrac{1}{\det M} \cdot \cfrac{1}{\sqrt{5}} \begin{bmatrix}
-1 & -2 \\ -2 & 1
\end{bmatrix}^T = \cfrac{1}{\det M} \cdot \cfrac{1}{\sqrt{5}} \begin{bmatrix}
-1 & -2 \\ -2 & 1
\end{bmatrix}"

Determinant of matrix "M" equals to:

"\det M = \cfrac{1}{\sqrt{5}} \cdot (-\cfrac{1}{\sqrt{5}}) - \cfrac{2}{\sqrt{5}} \cdot \cfrac{2}{\sqrt{5}} = -\cfrac{1}{5} - \cfrac{4}{5}=-1"

So, inverse matrix "M^{-1}" is equal to:

"M^{-1} = \cfrac{1}{\det M} \cdot \cfrac{1}{\sqrt{5}} \begin{bmatrix}
-1 & -2 \\ -2 & 1
\end{bmatrix} = \cfrac{1}{-1} \cdot \cfrac{1}{\sqrt{5}} \begin{bmatrix}
-1 & -2 \\ -2 & 1
\end{bmatrix} = \cfrac{1}{\sqrt{5}} \begin{bmatrix}
1 & 2 \\ 2 & -1
\end{bmatrix}"

So, matrix "D" equals to:

"D = M^{-1} A M = \cfrac{1}{\sqrt{5}} \begin{bmatrix}
1 & 2 \\ 2 & -1
\end{bmatrix} \cdot \begin{bmatrix}
8 & -2 \\ -2 & 5
\end{bmatrix} \cdot \cfrac{1}{\sqrt{5}} \begin{bmatrix}
1 & 2 \\ 2 & -1
\end{bmatrix} = \cfrac{1}{5} \cdot \begin{bmatrix}
1 & 2 \\ 2 & -1
\end{bmatrix} \cdot \begin{bmatrix}
8 & -2 \\ -2 & 5
\end{bmatrix} \cdot \begin{bmatrix}
1 & 2 \\ 2 & -1
\end{bmatrix} = \cfrac{1}{5} \begin{bmatrix}
4 & 8 \\ 18 & -9
\end{bmatrix} \cdot \begin{bmatrix}
1 & 2 \\ 2 & -1
\end{bmatrix} = \cfrac{1}{5} \cdot \begin{bmatrix}
20 & 0 \\
0 & 45
\end{bmatrix} = \begin{bmatrix}
4 & 0 \\
0 & 9
\end{bmatrix}"

So, the canonical form can be found as:

"D_{11} \hat{x}^2 + D_{22} \hat{y}^2 =4\hat{x}^2+9\hat{y}^2"

Answer: "4\hat{x}^2+9\hat{y}^2"